OpenStudy (marigirl):

solve 8 cos^3(x)=1 Thanks!

Nnesha (nnesha):

solve for x ? find exact value ?

OpenStudy (marigirl):

\[8 \cos^3(x)=1\] \[\cos^3(x)=\frac{ 1 }{ 8 }\] \[\cos(x)=\sqrt[3]{\frac{ 1 }{ 8 }}\]

Nnesha (nnesha):

looks correct :=

Nnesha (nnesha):

do you need help from there ??

Nnesha (nnesha):

\[\huge\rm \sqrt{\frac{ a }{ b }} = \frac{ \sqrt{a} }{ \sqrt{b} }\]

Nnesha (nnesha):

same with \[\large\rm \sqrt[2]{\frac{ 1 }{ 8 }} =\frac{ \sqrt[3]{1} }{ \sqrt[3]{8}}\]